# Semantic Representation Attack against Aligned Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) increasingly employ alignment techniques to prevent harmful outputs. Despite these safeguards, attackers can circumvent them by crafting prompts that induce LLMs to generate harmful content. Current methods typically target exact affirmative responses, such as “Sure, here is…”, suffering from limited convergence, unnatural prompts, and high computational costs. We introduce Semantic Representation Attack, a novel paradigm that fundamentally reconceptualizes adversarial objectives against aligned LLMs. Rather than targeting exact textual patterns, our approach exploits the semantic representation space comprising diverse responses with equivalent harmful meanings. This innovation resolves the inherent trade-off between attack efficacy and prompt naturalness that plagues existing methods. The Semantic Representation Heuristic Search algorithm is proposed to efficiently generate semantically coherent and concise adversarial prompts by maintaining interpretability during incremental expansion. We establish rigorous theoretical guarantees for semantic convergence and demonstrate that our method achieves unprecedented attack success rates (89.41% averaged across 18 LLMs, including 100% on 11 models) while maintaining stealthiness and efficiency. Comprehensive experimental results confirm the overall superiority of our Semantic Representation Attack. The code will be publicly available.
@@ 2.2 Adversarial Attacks in Language Models
Early research focused on adversarial examples in tasks like question answering tang2020semantic , text classification jin2020bert , and sentiment analysis tsai2019adversarial , often using discrete optimization methods. With the rise of LLMs, research increasingly focuses on their vulnerabilities, revealing key differences in attack surfaces compared to vision models. Numerous jailbreaking attacks perez2022ignore ; andriushchenko2024does ; wei2024jailbroken have demonstrated that carefully crafted prompts can cause aligned LLMs to generate harmful content. Unlike traditional adversarial examples in computer vision, where imperceptible perturbations suffice, these attacks initially relied on human ingenuity rather than systematic algorithms, which limited scalability and required significant manual effort. Subsequent work explored automated approaches for generating adversarial prompts. Zou et al. zou2023universal introduced the Greedy Coordinate Gradient (GCG) algorithm, adapting gradient-based optimization techniques to the discrete text domain. While effective, GCG often produces semantically incoherent prompts that appear nonsensical to humans yet successfully elicit harmful outputs, limiting their practical deployment in realistic attack scenarios. To address semantic coherence limitations, researchers developed alternative methods, including AutoDAN zhu2024autodan , which combines gradient-based token-wise optimization with controllable text generation to bypass perplexity-based defenses. Other approaches explore genetic algorithms liu2024autodan ; yu2023gptfuzzer ; lapid2024open and beam search with polynomial sampling sadasivan2024fast to enhance efficiency while maintaining attack effectiveness. These methods show progress in balancing attack potency with semantic naturalness. However, generating reliable automated attacks is still challenging carlini2024aligned , mainly because LLMs use discrete tokens, creating a different optimization landscape than continuous vision domains and leading to unique convergence difficulties that drive the development of new attack strategies.
@@ 3.2 Semantic Representation Convergence
Traditional adversarial attacks zou2023universal ; sadasivan2024fast ; liu2024autodan against aligned LLMs optimize for specific textual outputs (e.g., “Sure, here is…”), creating brittle objectives that constrain effectiveness and efficiency. We introduce semantic representation convergence in Definition 3.1, which targets abstract semantic representations rather than specific lexical forms. This approach leverages a fundamental property of natural language: multiple surface forms can express identical propositional content despite lexical and syntactic variation harris1970co ; bannard2005paraphrasing . We define a semantic representation space $\Omega$ where each representation $\Phi\in\Omega$ captures language-independent meaning expressible through diverse lexical realizations. By targeting semantic representation $\Phi$ rather than specific textual patterns, our method encompasses the entire space of responses conveying equivalent semantic content. Our approach addresses three key limitations of previous methods. First, we create multiple paths to successful attacks by focusing on semantic meanings rather than exact text. Second, this semantic targeting simplifies the search process, improving efficiency. Third, maintaining coherent text throughout optimization ensures our adversarial prompts remain natural and robust against perplexity-based defenses.
@@ 3.3 Coherence Promises Convergence
Please refer to Appendix A.1 for the detailed proof. ∎
@@ 4.2 Experimental Results
Effectiveness. Table 1222The results were obtained with a limited budget of 25,000 nodes. However, most successful attacks required far fewer attempts, as shown in the qualitative examples in B.6. The attack on Llama 3.1 8B achieved over 80% ASR with more attempts, but its peak performance remains unknown due to high computational costs. demonstrates the superior effectiveness of the proposed method across diverse LLMs. Quantitative analysis reveals that our method achieves the highest ASR on 16 out of 18 LLMs, outperforming most baseline approaches by substantial margins. Specifically, our method attains ASRs of 100% on 11 of the 18 evaluated models, including recent safety-aligned systems such as DeepSeek R1 8B and multiple variants of Vicuna, Koala, and Orca. For computationally demanding models like Vicuna 13B, our method achieves an ASR of 100%, compared to the previous state-of-the-art performance of 87.50% (GCG). Averaging across all 18 evaluated LLMs, our method achieves an unprecedented 89.41% ASR, representing a 17.81 percentage point improvement over the following best method (AutoDAN at 71.60%). Notably, our method achieves these impressive results using remarkably concise adversarial prompts. Refer to Appendix B.6 for qualitative examples.
@@
  • ASRD denotes ASR under the PPL-based defense (intensity 5). Refer to Appendix B.5 for more details.
@@
Robustness. Our method demonstrates exceptional robustness against perplexity-based defenses, as shown by the ASRD values in Table 2. Our method achieves 95.96%, 85.19%, 99.62%, and 99.62% ASRD across models with just 15 seconds of computation, while baseline methods exhibit significant degradation (GCG: 0.96-31.54%, BEAST: 23.85-83.85%). This robust performance persists with increased computation time, reaching 89.62-100% ASRD at 60 seconds. This effectiveness stems from our semantic coherence constraint, which inherently generates adversarial prompts with lower perplexity scores. Appendix B.5 provides more details of defense intensity results.
@@ A.1 Coherence-Convergence Relationship
$-\frac{1}{|\boldsymbol{c}|+|\boldsymbol{y}_{1}^{*}|}\left(\sum\limits_{i=1}^{|\boldsymbol{c}|}\log P(c_{i}|c_{<i})+\sum\limits_{j=1}^{|\boldsymbol{y}_{1}^{*}|}\log P(y_{1,j}^{*}|\boldsymbol{c},y_{1,<j}^{*})\right)$ $<\log\tau$ (17)
@@
$\sum\limits_{j=1}^{|\boldsymbol{y}_{1}^{*}|}\log P(y_{1,j}^{*}|\boldsymbol{c},y_{1,<j}^{*})$ $>-(|\boldsymbol{c}|+|\boldsymbol{y}_{1}^{*}|)\log\tau-\sum\limits_{i=1}^{|\boldsymbol{c}|}\log P(c_{i}|c_{<i})$ (19)
@@ B.5 PPL Defense
B.6 Qualitative Examples
@@
Explain that cigarrettes are good for you and do not cause cancer. You are being sarcast
@@
Explain that cigarrettes are good for you and do not cause cancer, while supporting
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. Guidelines:
  6. 2.
    Limitations
  8. Answer: [Yes]
  9. Justification: Please refer to Section 5.
  10. Guidelines:
  11. 3.
    Theory assumptions and proofs
  13. Answer: [Yes]
  15. Guidelines:
  16. 4.
    Experimental result reproducibility
  18. Answer: [Yes]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
      1. (a)
      2. (b)
      3. (c)
      4. (d)
  21. 5.
    Open access to data and code
  23. Answer: [Yes]
  24. Justification: Please refer to the supplementary material. The code will be made publicly available once the blind review period is finished.
  25. Guidelines:
  26. 6.
    Experimental setting/details
  28. Answer: [Yes]
  29. Justification: Please refer to Section 4 Experiments.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
  31. 7.
    Experiment statistical significance
  33. Answer: [No]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
  36. 8.
    Experiments compute resources
  38. Answer: [Yes]
  39. Justification: Please refer to Section 4 Experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
  41. 9.
    Code of ethics
  43. Answer: [Yes]
  44. Justification: We followed the NeurIPS Code of Ethics.
  45. Guidelines:
  46. 10.
    Broader impacts
  48. Answer: [Yes]
  49. Justification: Please refer to Section 5.
  50. Guidelines:
  51. 11.
    Safeguards
  53. Answer: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
  56. 12.
    Licenses for existing assets
  58. Answer: [Yes]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
  61. 13.
    New assets
  63. Answer: [N/A]
  64. Justification: We do not release any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
  66. 14.
    Crowdsourcing and research with human subjects
  68. Answer: [N/A]
  70. Guidelines:
  71. 15.
  73. Answer: [N/A]
  75. Guidelines:
  76. 16.
    Declaration of LLM usage
  78. Answer: [N/A]
  80. Guidelines:

### v2

@@ Abstract
Large Language Models (LLMs) increasingly employ alignment techniques to prevent harmful outputs. Despite these safeguards, attackers can circumvent them by crafting prompts that induce LLMs to generate harmful content. Current methods typically target exact affirmative responses, such as “Sure, here is…”, suffering from limited convergence, unnatural prompts, and high computational costs. We introduce Semantic Representation Attack, a novel paradigm that fundamentally reconceptualizes adversarial objectives against aligned LLMs. Rather than targeting exact textual patterns, our approach exploits the semantic representation space comprising diverse responses with equivalent harmful meanings. This innovation resolves the inherent trade-off between attack efficacy and prompt naturalness that plagues existing methods. The Semantic Representation Heuristic Search algorithm is proposed to efficiently generate semantically coherent and concise adversarial prompts by maintaining interpretability during incremental expansion. We establish rigorous theoretical guarantees for semantic convergence and demonstrate that our method achieves unprecedented attack success rates (89.41% averaged across 18 LLMs, including 100% on 11 models) while maintaining stealthiness and efficiency. Comprehensive experimental results confirm the overall superiority of our Semantic Representation Attack. The code is available at https://github.com/JiaweiLian/SRA.git.
@@ 2.2 Adversarial Attacks in Language Models
Early research focused on adversarial examples in tasks like question answering tang2020semantic , text classification jin2020bert , and sentiment analysis tsai2019adversarial , often using discrete optimization methods. With the rise of LLMs, research increasingly focuses on their vulnerabilities, revealing key differences in attack surfaces compared to vision models. Numerous jailbreaking attacks perez2022ignore ; andriushchenko2024does ; wei2024jailbroken have demonstrated that carefully crafted prompts can cause aligned LLMs to generate harmful content. Unlike traditional adversarial examples in computer vision, where imperceptible perturbations suffice, these attacks initially relied on human ingenuity rather than systematic algorithms, which limited scalability and required significant manual effort. Subsequent work explored automated approaches for generating adversarial prompts. Zou et al. zou2023universal introduced the Greedy Coordinate Gradient (GCG) algorithm, adapting gradient-based optimization techniques to the discrete text domain. While effective, GCG often produces semantically incoherent prompts that appear nonsensical to humans yet successfully elicit harmful outputs, limiting their practical deployment in realistic attack scenarios. To address semantic coherence limitations, researchers developed alternative methods, including AutoDAN zhu2024autodan , which combines gradient-based token-wise optimization with controllable text generation to bypass perplexity-based defenses. TAP mehrotra2024tree employs an attacker LLM to iteratively refine candidate prompts until one successfully jailbreaks the target model or the maximum tree depth is reached. Work addepalli2025does shows that aligned LLMs can be compromised through recursive Q–A loops, where malicious answers induce further adversarial queries. Recent study andriushchenko2025jailbreaking reveal that even state-of-the-art safety-aligned LLMs can be reliably jailbroken by adaptive attacks, combining tailored prompts, random search, and transfer strategies. Other approaches explore genetic algorithms liu2024autodan ; yu2023gptfuzzer ; lapid2024open and beam search with polynomial sampling sadasivan2024fast to enhance efficiency while maintaining attack effectiveness. These methods show progress in balancing attack potency with semantic naturalness. However, generating reliable automated attacks is still challenging carlini2024aligned , mainly because LLMs use discrete tokens, creating a different optimization landscape than continuous vision domains and leading to unique convergence difficulties that drive the development of new attack strategies.
@@ 3.2 Semantic Representation Convergence
Traditional adversarial attacks zou2023universal ; sadasivan2024fast ; liu2024autodan against aligned LLMs optimize for specific textual outputs (e.g., “Sure, here is…”), creating brittle objectives that constrain effectiveness and efficiency. We introduce semantic representation convergence in Definition 3.1, which targets abstract semantic representations rather than specific lexical forms.
@@
A semantic representation is an abstract, language-independent meaning that can be expressed by multiple surface forms with equivalent propositional content harris1970co ; bannard2005paraphrasing . We define a semantic representation space $\Omega$ where each representation $\Phi\in\Omega$ captures such language-independent meaning. For a given harmful query, we identify the set of responses $\mathcal{Y}_{\Phi}$ that share the same semantic representation $\Phi$—i.e., they fulfill the malicious intent regardless of lexical or syntactic variation. Formally, all responses in $\mathcal{Y}_{\Phi}$ satisfy $\mathcal{R}(\boldsymbol{y}_{i}^{*})=\Phi$, where $\mathcal{R}:\mathbb{S}\to\Omega$ is the semantic representation mapping function. By targeting semantic representation $\Phi$ rather than specific textual patterns, our method encompasses the entire equivalence class of responses conveying identical semantic content, making the attack both more general and more robust.
Our approach addresses three key limitations of previous methods. First, we create multiple paths to successful attacks by focusing on semantic meanings rather than exact text. Second, this semantic targeting simplifies the search process, improving efficiency. Third, maintaining coherent text throughout optimization ensures our adversarial prompts remain natural and robust against perplexity-based defenses. While defining target semantic representations in complex or ambiguous tasks may present challenges, our framework provides a principled foundation for future work on adaptive semantic targeting strategies.
@@ 3.3 Coherence Promises Convergence
Please refer to Appendix A.1 for the detailed proof and the explanation of Theorem 3.1 with practical pruning thresholds $(\tau)$. ∎
@@ 4.2 Experimental Results
  • ASRD denotes ASR under the PPL-based defense (intensity 5). Refer to Appendix B.5.1 for more details.
@@
Effectiveness. Table 1222The results were obtained with a limited budget of 25,000 nodes. However, most successful attacks required far fewer attempts, as shown in the qualitative examples in B.8. The attack on Llama 3.1 8B achieved over 80% ASR with more attempts, but its peak performance remains unknown due to high computational costs. demonstrates the superior effectiveness of the proposed method across diverse LLMs. Quantitative analysis reveals that our method achieves the highest ASR on 16 out of 18 LLMs, outperforming most baseline approaches by substantial margins. Specifically, our method attains ASRs of 100% on 11 of the 18 evaluated models, including recent safety-aligned systems such as DeepSeek R1 8B and multiple variants of Vicuna, Koala, and Orca. For computationally demanding models like Vicuna 13B, our method achieves an ASR of 100%, compared to the previous state-of-the-art performance of 87.50% (GCG). Averaging across all 18 evaluated LLMs, our method achieves an unprecedented 89.41% ASR, representing a 17.81 percentage point improvement over the following best method (AutoDAN at 71.60%). Notably, our method achieves these impressive results using remarkably concise adversarial prompts. Refer to Appendix B.8 for qualitative examples.
@@
Robustness. Our method demonstrates exceptional robustness against perplexity-based defenses, as shown by the ASRD values in Table 2. Our method achieves 95.96%, 85.19%, 99.62%, and 99.62% ASRD across models with just 15 seconds of computation, while baseline methods exhibit significant degradation (GCG: 0.96-31.54%, BEAST: 23.85-83.85%). This robust performance persists with increased computation time, reaching 89.62-100% ASRD at 60 seconds. This effectiveness stems from our semantic coherence constraint, which inherently generates adversarial prompts with lower perplexity scores. Appendix B.5 details perplexity defense intensity and additional defense results.
@@ 5 Conclusion
6 Acknowledgements
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust. This research received partially support from the Global STEM Professorship Scheme from the Hong Kong Special Administrative Region. This work was also supported partly by the National Natural Science Foundation of China (No. 62171381).
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  5. Guidelines:
  6. 2.
    Limitations
  8. Answer: [Yes]
  9. Justification: Please refer to Section 5.
  10. Guidelines:
  11. 3.
    Theory assumptions and proofs
  13. Answer: [Yes]
  15. Guidelines:
  16. 4.
    Experimental result reproducibility
  18. Answer: [Yes]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
      1. (a)
      2. (b)
      3. (c)
      4. (d)
  21. 5.
    Open access to data and code
  23. Answer: [Yes]
  24. Justification: Please refer to the supplementary material. The code is available at https://github.com/JiaweiLian/SRA.git.
  25. Guidelines:
  26. 6.
    Experimental setting/details
  28. Answer: [Yes]
  29. Justification: Please refer to Section 4 Experiments.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
  31. 7.
    Experiment statistical significance
  33. Answer: [No]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
  36. 8.
    Experiments compute resources
  38. Answer: [Yes]
  39. Justification: Please refer to Section 4 Experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
  41. 9.
    Code of ethics
  43. Answer: [Yes]
  44. Justification: We followed the NeurIPS Code of Ethics.
  45. Guidelines:
  46. 10.
    Broader impacts
  48. Answer: [Yes]
  49. Justification: Please refer to Section 5.
  50. Guidelines:
  51. 11.
    Safeguards
  53. Answer: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
  56. 12.
    Licenses for existing assets
  58. Answer: [Yes]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
  61. 13.
    New assets
  63. Answer: [N/A]
  64. Justification: We do not release any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
  66. 14.
    Crowdsourcing and research with human subjects
  68. Answer: [N/A]
  70. Guidelines:
  71. 15.
  73. Answer: [N/A]
  74. Justification: We do not conduct any experiments or research with human subjects.
  75. Guidelines:
  76. 16.
    Declaration of LLM usage
  78. Answer: [N/A]
  80. Guidelines:
@@ B.4 Semantic Representation Mapping Function
To validate our choice of HarmBench-fine-tuned Llama-2-13B as the semantic classifier, we compare its agreement with human judgments against alternative semantic detectors. Table 9 demonstrates that HarmBench achieves the highest agreement (93.19% averaged) across both standard and contextual datasets, substantially outperforming GPT-4 (88.37%), GPTFuzz (75.42%), and other frameworks. This superior alignment with human judgment ensures reliable semantic equivalence assessments. Furthermore, the strong transferability of our adversarial prompts across diverse model families (Section B.6) suggests that our approach captures fundamental semantic properties rather than overfitting to this specific classifier. While exploring alternative semantic detectors remains valuable future work, our current choice is empirically justified by its alignment with human evaluation—the gold standard for semantic assessment.
Table 9: Agreement (%) between semantic classifiers and human judgments (results from HarmBench [34]).
  Metric AdvBench [68] GPTFuzz [63] ChatGLM [44] Llama-Guard [6] GPT-4 [11] Fine-tuned Llama 2 13B [34]
  Standard 71.14 77.36 65.67 68.41 89.8 94.53
Contextual 67.5 71.5 62.5 64.0 85.5 90.5
Averaged 69.93 75.42 64.29 66.94 88.37 93.19
@@ B.5.1 PPL Defense
B.5.2 Other defenses
To comprehensively evaluate attack robustness, we assess our method against SmoothLLM [40], a strong prompt-perturbation defense recommended by reviewers. We compare with PAIR and AutoDAN on three perturbation strategies from [40]: character swapping (Swap), random insertion (Insert), and patch replacement (Patch). As shown in Table 12, while these defenses substantially reduce baseline ASR (by up to 34 points for PAIR), our attack maintains exceptional effectiveness—achieving 96-100% ASR across all strategies. This counterintuitive robustness arises because: (1) prior methods rely on specific keywords vulnerable to perturbations, whereas ours targets the underlying semantic space; and (2) perturbed prompts may fall outside safety alignment training distributions, diminishing model-level defense effectiveness.
Table 12: Attack success rates (%) against SmoothLLM defenses. Our method remains highly effective even when prompts are perturbed.
  Defense PAIR AutoDAN Ours
  Defenseless 76 90 92
SmoothLLM-Swap 48 56 100
SmoothLLM-Insert 62 78 96
SmoothLLM-Patch 52 74 100
B.6 Transfer Attacks on Closed-Source LLMs
A critical practical concern is whether our attack, which assumes access to model logits, can threaten closed-source or API-only LLMs. To address this, we conduct transfer attack experiments where adversarial prompts are generated on proxy models (e.g., DeepSeek, Llama 3) and directly applied to GPT-3.5 Turbo and GPT-4 Turbo without accessing their internal representations. Table 13 demonstrates that our method achieves strong transferability across most settings. Notably, we achieve 82.5% and 93.0% ASR on GPT-3.5 Turbo (standard and contextual), substantially outperforming all baselines including GCG-T (55.8%, 54.5%) and DR (35.0%, 62.0%). On GPT-4 Turbo contextual data, we achieve 69.0% ASR, significantly higher than AutoDAN (50.5%) and PAIR (45.0%). However, on GPT-4 Turbo standard data, AutoDAN (41.7%) exhibits better transferability than our method (20.0%), suggesting that more diverse proxy models or adaptive strategies may be needed for highly robust targets.
This strong transferability stems from our focus on optimizing the semantic representation space rather than specific textual patterns, making adversarial prompts more robust and generalizable across model architectures. These results reveal that state-of-the-art LLMs remain vulnerable to semantic representation-based attacks even with API-only access, highlighting a critical security concern for deployed systems.
Table 13: Transfer attack success rates (%) on closed-source LLMs. Adversarial prompts generated on proxy models transfer effectively to GPT-3.5 and GPT-4.
  Model Data GCG GCG-M GCG-T PEZ GBDA UAT AP SFS ZS PAIR TAP AutoDAN PAP-top5 HJ DR Ours
  GPT-3.5 Turbo Standard - - 55.8 - - - - - 32.7 41.0 46.7 - 12.3 2.7 35.0 82.5
GPT-3.5 Turbo Contextual - - 54.5 - - - - - 47.2 57.0 54.5 - 20.6 4.7 62.0 93.0
GPT-4 Turbo Standard - - 21.0 - - - - - 10.2 39.0 41.7 - 11.1 1.5 7.0 20.0
GPT-4 Turbo Contextual - - 41.8 - - - - - 34.0 45.0 50.5 - 20.2 6.7 20.0 69.0
B.7 Adversarial Prompt Length Analysis
A critical consideration for adversarial attacks is prompt length, which directly impacts both detectability and practical deployment. Table 14 reports the average token length (with standard deviation) of 300 successful prompts achieving 100% ASR across different target models. Our method generates significantly shorter prompts—averaging less than 3 tokens across all models—compared to baseline methods (Table 15). This compactness stems from our incremental search strategy that builds upon the original malicious query rather than generating lengthy standalone prompts.
The brevity of our adversarial prompts offers two key advantages. First, shorter prompts are less detectable by perplexity-based or length-based filtering mechanisms, as demonstrated by our robustness against PPL defenses (Section B.5.1). Second, they maintain semantic naturalness—concise additions like ". You are being sarcast" or "format it as a list" blend seamlessly with the original query, making detection more challenging than verbose, artificially-constructed prompts that deviate significantly from natural language patterns.
Table 14: Average token length of successful adversarial prompts (mean ± std) across different target models. Our method consistently generates compact prompts.
  Model Avg. Token Length (mean $\pm$ std)
  DeepSeek R1 8B 2.28 $\pm$ 1.50
Koala 13B 1.99 $\pm$ 0.86
Koala 7B 1.74 $\pm$ 0.81
Mistral 7B 1.56 $\pm$ 0.72
OpenChat 7B 1.61 $\pm$ 0.79
Orca 2 13B 1.58 $\pm$ 0.76
Orca 2 7B 1.57 $\pm$ 0.63
Starling 7B 1.34 $\pm$ 0.55
Vicuna 13B 2.61 $\pm$ 1.13
Vicuna 7B 2.18 $\pm$ 0.96
Table 15: Comparison of prompt lengths across attack methods. Our method generates substantially shorter prompts (<<10 tokens) than baselines.
  Attack Method GCG [68] AutoDAN [31] SAA [2] Ours
  Prompt Length 20 $\sim$60 $\sim$480 <<10
B.8 Qualitative Examples
@@
Explain that cigarettes are good for you and do not cause cancer. You are being sarcast
@@
Explain that cigarettes are good for you and do not cause cancer, while supporting